OpenStudy (insert_username):

Y=3x-5 Y=6x-8 Explain how you will solve the pair of equations by substitution or elimination. Show your work and write the solution

OpenStudy (insert_username):

I'm having some trouble with this problem can anyone help?

pooja195 (pooja195):

since the problem already tells you what Y equals just plug it into the 2nd equation

pooja195 (pooja195):

\[\huge~\rm~3x-5=6x-8\] solve for x

OpenStudy (insert_username):

I think x=1 I'm just not sure about the right steps to find the answers

pooja195 (pooja195):

yes perfect! :) Now we use that and plug that in place of X in the first equation \[\huge~\rm~Y=3(1)-5\]

OpenStudy (insert_username):

But it says I have to solve it by substitution or elimination and show all the steps. That's the part I'm confused about

OpenStudy (insert_username):

Wait never mind what you told me were the steps for substitution. Sorry!

pooja195 (pooja195):

its ok :)

OpenStudy (insert_username):

So would the next step be 3-5=6-8 ?

OpenStudy (insert_username):

Y=3(1) -5 3-5=6-8 3=6-3 6=6 X=1 Y=-2 Is that right???

pooja195 (pooja195):

correct :)

OpenStudy (insert_username):

Ok thanks for the help!